# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ 4.3 Frequentist Coverage
As described earlier, for each $(n,p)$, we first consider $100$ randomly chosen variables and hold them fixed across $R=100$ replicates of the data. Next, we obtain the average coverage and interval width of $95\%$ posterior/pseudo-posterior credible intervals corresponding to the relevant $100\times 100$ submatrix of $\Psi_{0}$. For all the simulation instances, we construct the entries $b_{uv}$ as described in Section 2.2, and then proceed to set the coverage-correction factor $\rho$ to be the average of $\{b_{uv}:1\leq u\leq v\leq p\}$ for coverage-correction. The results, provided in Table 2, suggest that CC-FABLE provides superior entrywise coverage when compared to MGSP, with comparable interval width. In all the cases, the average coverage of entrywise intervals obtained from MGSP falls short of the nominal value $0.95$, while CC-FABLE meets the nominal coverage. Furthermore, the average entrywise coverage obtained from MGSP decreases when the number of dimensions increases for a fixed sample size, while the results of CC-FABLE are not affected.
Table 2: Comparison of average frequentist coverages and average interval widths across entrywise pseudo-credible intervals obtained from CC-FABLE and MGSP.
Cases Coverage Coverage Width Width
$(n,p)$ (CC-FABLE) (MGSP) (CC-FABLE) (MGSP)
$n=500,p=1000$ 0.95 0.84 0.48 0.48
$n=1000,p=1000$ 0.95 0.72 0.35 0.31
$n=500,p=5000$ 0.95 0.81 0.50 0.46
$n=1000,p=5000$ 0.95 0.69 0.35 0.31
@@ 4.4 Computational Efficiency
For the purposes of computing order complexity, we assume $k$ is fixed and $n,p$ vary. We first consider the order complexity when generating $N_{0}$ pseudo-posterior samples of the covariance matrix. The first stage SVD of the $n\times p$ matrix $\mathbf{Y}$ has $\mathcal{O}(np\min(n,p))$ complexity (Vasudevan and Ramakrishna,, 2017), which needs to be only computed once. Computing the $p$ hyperparameters $\mu_{j}$ and $\delta_{j}^{2}$ for $j=1,\ldots,p$ has $\mathcal{O}(np)$ complexity, while generating each Monte Carlo sample of $({\Lambda}_{C},{\Sigma}_{C})$ for $j=1,\ldots,p$ has $\mathcal{O}(p)$ complexity since $({\lambda}_{C,j},{\sigma}_{C,j}^{2})$ are independently generated for $j=1,\ldots,p$. Lastly, obtaining $\Psi_{C}=\Lambda_{C}\Lambda_{C}^{\top}+\Sigma_{C}$ has $\mathcal{O}(p^{2})$ complexity for each sample, leading to $\mathcal{O}(p^{2}N_{0})$ complexity to obtain $N_{0}$ samples. Thus, obtaining $N_{0}$ pseudo-posterior samples of the covariance matrix after coverage-correction has $\mathcal{O}\{p^{2}N_{0}+pN_{0}+np\min(n,p)\}=\mathcal{O}\{p^{2}N_{0}+np\min(n,p)\}$ complexity. Next, we consider the order complexity to obtain the pseudo-posterior mean of the covariance. As before, we would require computing the SVD with $\mathcal{O}\{np\min(n,p)\}$ complexity, while computing the hyperparameters $(\mu_{j},\delta_{j}^{2})$ for $j=1,\ldots,p$ has $\mathcal{O}(np)$ complexity. Obtaining the matrix $G=G_{0}G_{0}^{\top}$ from $G_{0}$ has $\mathcal{O}(p^{2})$ complexity. Thus, computing the pseudo-posterior mean of $\Psi$ has asymptotic complexity $\mathcal{O}\{p^{2}+np\min(n,p)\}$.
The runtime for both sampling and pseudo-posterior mean computation may be substantially improved by considering computationally efficient implementations of the SVD step such as the randomized SVD (Halko et al.,, 2011) when either $n$ or $p$ is very large. Additionally, since the CC-FABLE algorithm employs embarrassingly parallel sampling, the computational runtime would be further improved by parallelization of the computations across cores.
In Figure 1, we compare the three methods CC-FABLE, MGSP, and ROTATE in terms of obtaining (pseudo) posterior samples and point estimates of the covariance matrix. We consider the same data-generating scenario as in Section 4.1. We fix $n=500$ and vary $p$ from $500$ to $5000$ in increments of $500$, obtaining runtime measurements across $R=20$ independent replicates for each choice of $p$. As before, we consider $3000$ MCMC iterations for MGSP and discard the first $1000$ iterates as burn-in. The experiments were carried out on an M1 MacBook Pro with 32 GB of RAM (random-access memory). We compare CC-FABLE with MGSP when obtaining samples of the factor loadings and the error variances, and with ROTATE when obtaining a point estimate of the covariance matrix. In both cases, CC-FABLE is faster compared to its competitors. When obtaining samples, CC-FABLE is about $150-200$ times faster than MGSP, with this ratio remaining fairly constant as $p$ increases. When obtaining a point estimator of the covariance, CC-FABLE is at least twice as fast as ROTATE.
Refer to caption Refer to caption
Figure 1: Plots comparing the runtimes in seconds (in $\log_{10}$ scale) of CC-FABLE, MGSP, and ROTATE approaches. The first plot compares the runtimes of CC-FABLE and MGSP when obtaining (pseudo) posterior samples of the covariance matrix. The second plot compares runtimes of CC-FABLE and ROTATE when obtaining a point estimate of the covariance matrix.
@@ 5 Application
In this Section, we apply our method to identify associations between expressed genes in a high-throughput sequencing immune cell dataset. Identifying gene expression associations in varying immune cell populations and among multiple lineages is critical to understanding the inflammatory response to infectious and chronic diseases (Ota et al.,, 2021; Saini et al.,, 2022). This enables subsequent detection of corresponding cellular and molecular pathways underlying pathological conditions. For this application, we consider the GSE109125 dataset, which contains RNAseq data from 127 highly purified immune cell lineages including adaptive and innate lymphocytes, myeloid cells, mast cells, and neutrophils. The lineages along with the associated differentiation cascades were profiled using the ImmGen ULI pipeline (Yoshida et al.,, 2019).
@@
We first log-transformed raw data counts $R_{ij}$ for the $i$th cell and $j$th gene as $y^{*}_{ij}=\log_{2}(R_{ij}+1)$. Next, we proceed to center the data for analysis before fitting the latent factor model. Finally, we use the genefilter package in R (Gentleman et al.,, 2023) to filter relevant genes. In this case, we filtered genes according to their variance and considered the top $10\%$ of genes with the highest variances. There were no cells with missing entries. After carrying out appropriate pre-processing of the dataset, we obtain expression data $\mathbf{Y}\in\mathbb{R}^{n\times p}$ for $p=5300$ genes measured on $n=205$ cells.
@@
We compare results from applying the three methods CC-FABLE, MGSP, and ROTATE on the data. The fitted log-likelihoods using the whole data $\mathbf{Y}$ are $-490006.4$, $-503574$, and $-728959.9$ for CC-FABLE, MGSP, and ROTATE, respectively. This suggests that CC-FABLE fits the data the best among the three methods. The estimated rank when using the $\mbox{JIC}(k)$ criterion as highlighted in Section 2.3 is $\widehat{k}=30$. As a posterior predictive check, we investigated the average coverage of entrywise predictive intervals across all the entries of $\mathbf{Y}$. When using CC-FABLE, the average predictive coverage is $0.97$. To further investigate model fit, we also looked into the proportion of variance explained for each gene using CC-FABLE. The average proportion of variance explained across variables when using CC-FABLE is $0.83$, indicating that the factor model is a good fit.
We now focus on estimating a $100\times 100$ submatrix of the $5300\times 5300$ covariance matrix, corresponding to the $100$ genes with the highest variability. Suppose the indices of these genes are $\mathcal{I}_{0}\subset\{1,2,\ldots,p\}$. We assess estimation performance using out-of-sample fit on a randomly chosen $50\times 100$ submatrix $\mathbf{Y}_{T}$ of $\mathbf{Y}$, corresponding to expression data on the $100$ genes in $\mathcal{I}_{0}$ for $50$ cells. To estimate the covariance of genes in $\mathcal{I}_{0}$, it is typical to only use expression data for those genes. Alternatively, we infer the covariance of these genes jointly with additional genes with indices in the set $\mathcal{A}(p_{S})$ of length $p_{S}>0$ such that $\mathcal{A}(p_{S})\cap\mathcal{I}_{0}=\phi$. We vary $p_{S}\in\{100j:j=1,\ldots,10\}\cup\{2000,4000\}$ and extract the relevant $100\times 100$ covariance $\widehat{\Psi}(p_{S})$ of the genes with indices in $\mathcal{I}_{0}$. We then use $\widehat{\Psi}(p_{S})$ to evaluate the out-of-sample log-likelihood (OOSLL) on the held-out test set $\mathbf{Y}_{T}$. We carry out this analysis for $R=10$ independent replicates of the train-test split and compare the methods CC-FABLE, MGSP, and ROTATE on the basis of average OOSLL. The results are illustrated in Figure 2.
In Figure 2, it is clear that both CC-FABLE and MGSP exhibit superior performance to ROTATE. Both MGSP and ROTATE exhibit an essentially monotonically decreasing trend as extra genes are added, so including additional genes does not improve inferences on the genes of interest in using these methods. In sharp contrast, CC-FABLE has rapidly improved performance initially as extra genes are added, exhibiting a clear blessing of dimensionality. After a few hundred genes, the gain levels off, and eventually there is a modest decline, perhaps due to the need to add additional factors not related to the genes of interest when large numbers of additional genes are added.
The proposed method thus shows promising results in this high-dimensional gene expression data application and compares favourably to the two state-of-the-art gold standard approaches to Bayesian factor analysis. The results obtained from CC-FABLE come at a fraction of the computational budget compared to MGSP due to its embarrassingly parallel sampling scheme requiring no MCMC. When fitted on the full dataset on an M1 MacBook Pro with $32$ GB of RAM, CC-FABLE only took $1.1$ seconds to compute the pseudo-posterior mean while running MGSP with $3000$ MCMC and $1000$ burn-in iterates took $\sim 27$ minutes. This translates to a speed-up of close to $1600$ times when using CC-FABLE.
@@
Figure 2: Figure showing log-likelihood of the test set (y-axis) averaged over $R=10$ replicates of the train-test split, when using additional genes (x-axis) to estimate the covariance of the relevant set of $100$ variables, using CC-FABLE, MGSP, and ROTATE. The training set has $155$ cells and the test set has $50$ cells for each of the $R=10$ train-test split replicates.
@@ 6 Discussion
In this paper, we developed a computationally scalable approach to fit high-dimensional Bayesian latent factor models with Gaussian data. Leveraging on a blessing of dimensions, the proposed FABLE approach completely bypasses the need to carry out MCMC to provide pseudo-posterior samples of the covariance matrix for inference. By avoiding MCMC altogether, our approach averts the pitfalls of MCMC in higher dimensions, such as slow mixing and unsatisfactory convergence. Due to its embarrassingly parallel nature, the FABLE approach leads to immense computational benefits over current approaches dependent on MCMC. We developed a coverage-correction modification to FABLE called CC-FABLE with provable guarantees on asymptotic frequentist coverage of its pseudo-credible intervals. The CC-FABLE pseudo-posterior also enjoys desirable theoretical properties such as consistency and compares favorably to state-of-the-art approaches to Bayesian factor analysis from the viewpoint of estimation error, uncertainty quantification, and computational efficiency in our numerical experiments and gene expression data application.
This work naturally leads to several important directions for future research. The first direction is to model non-Gaussian data, such as binary or count data. Such data are routinely collected in ecological studies. For example, $y_{ij}$ for $i=1,\ldots,n$ and $j=1,\ldots,p$ could indicate the presence/absence or abundance of species $j$ in sample $i$. For presence/absence data, let $\mbox{P}(y_{ij}=1\mid\alpha_{0},\eta_{i},\lambda_{j})=\mbox{F}\left(\alpha_{0}+\eta_{i}^{\top}\lambda_{j}\right),$ with $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$. Here, $F$ is a link function, such as the probit or the logistic link. The observations $y_{ij}$ are mutually independent conditional on the latent factors $\eta_{i}$, but integrating them out induces dependence between $y_{i1},\ldots,y_{ip}.$ Bayesian inference in such models typically proceeds via a data augmentation Gibbs sampler by introducing auxiliary continuous variables $z_{ij}$ and alternating between sampling $\eta_{i},\lambda_{j},$ and $z_{ij}$ for $i=1,\ldots,n$ and $j=1,\ldots,p$ given the data $\mathbf{Y}=(y_{ij})$ using MCMC (Albert and Chib,, 1993; Polson et al.,, 2013). However, in our motivating ecological settings involving insects and fungi, $p$ is on the order of 10,000100,$0-100,00$ 100 , 000, rendering current MCMC-based Bayesian factor analytic approaches computationally infeasible. To extend FABLE to this setting, we can pre-estimate the latent factors via an SVD or an appropriate generalized linear model (GLM) variant, and then fit the resulting Bayesian GLMs in parallel for each species, potentially relying on Laplace approximations.
@@
An additional important direction is to modify the FABLE inferential framework to allow more intricate hierarchical modeling. For example in Bayesian factor analysis, it is often of interest to (1) include covariate effects, (2) use more elaborate priors on the loadings to favor sparsity and adaptive selection of the number of factors, (3) consider more flexible latent factor distributions, and/or (4) allow nonlinear measurement structure. In exploratory analyses, we have observed good performance for ad hoc FABLE modifications to include covariates and for simulated data having non-Gaussian latent factor distributions. Problem (2) can potentially be addressed in the second stage via MCMC algorithms implemented in parallel for the different sparse regressions after inferring the latent factors using a sparse SVD in a first stage. Problem (4) may necessitate non-linear dimensionality reduction in the first stage. Obtaining concrete implementations with theory support for problems (1)-(4) is of future interest.
@@
The current work also hints at further investigation into the theoretical properties of Bayesian factor models. Although there are existing results focusing on posterior consistency with carefully chosen shrinkage priors on the factor loadings, these are primarily in the case where the idiosyncratic model errors are Gaussian. It would thus be interesting to explore how the blessing of dimensions influences both posterior consistency and uncertainty quantification in different settings, such as when the data are non-Gaussian. Furthermore, there remains a clear need to extend existing theoretical results on Bayesian factor models to results on uncertainty quantification of the covariance matrix.

### v3

@@ 4.3 Frequentist Coverage
Table 3: Comparison of frequentist coverages and interval widths across entrywise credible intervals obtained from FABLE and MGSP. The “Mean” and “Range” columns show the mean and $2.5\%-97.5\%$ quantiles across replicates, respectively.
Method FABLE MGSP
Metric Coverage Width Coverage Width
$(n,p)$ Mean Range Mean Range Mean Range Mean Range
$(500,1000)$ 0.95 0.94 – 0.97 0.46 0.45 – 0.47 0.85 0.83 – 0.86 0.45 0.44 – 0.46
$(1000,1000)$ 0.95 0.94 – 0.96 0.32 0.32 – 0.33 0.80 0.79 – 0.81 0.32 0.31 – 0.33
$(500,5000)$ 0.96 0.94 – 0.97 0.48 0.46 – 0.49 0.83 0.81 – 0.84 0.44 0.43 – 0.45
$(1000,5000)$ 0.95 0.94 – 0.97 0.34 0.33 – 0.34 0.77 0.76 – 0.78 0.31 0.31 – 0.32
As described earlier, for each $(n,p)$, we first consider $100$ randomly chosen variables and hold them fixed across $R=100$ replicates of the data. Next, we obtain the average coverage and interval width of $95\%$ posterior credible intervals corresponding to the relevant $100\times 100$ submatrix of $\Psi_{0}$. For all the simulation instances, we set the coverage-correction factor $\rho$ as recommended in Section 2.3. The results, provided in Table 3, suggest that FABLE provides superior entrywise coverage when compared to MGSP, with comparable interval width. In all the cases, the average coverage of entrywise intervals obtained from MGSP falls short of the nominal value $0.95$, while FABLE meets the nominal coverage. Furthermore, the average entrywise coverage obtained from MGSP decreases when the number of dimensions increases for a fixed sample size, while the results of FABLE are not affected.
@@ 4.4 Computational Efficiency
[…]
The runtime for both sampling and the FABLE-posterior mean computation can be substantially improved by computationally efficient implementations of the SVD step, such as the randomized SVD (Halko et al.,, 2011) when either $n$ or $p$ is very large. Additionally, since FABLE employs embarrassingly parallel sampling, its computational runtime would be further improved by parallelizing the computations across cores.
In Figure 1, we compare FABLE with the other competitors in terms of obtaining posterior samples and point estimates of the covariance matrix. We consider the same data-generating scenario as in Section 4.1 with $\pi_{0}=0.5$. We fix $n=500$ and vary $p$ from $500$ to $4000$ in increments of $500$, obtaining runtime measurements across $R=20$ independent replicates for each choice of $p$. As before, we consider $3000$ MCMC iterations for MGSP and discard the first $1000$ iterates as burn-in. The experiments were carried out on an M1 MacBook Pro with 32 GB of RAM (random-access memory). We compare FABLE with MGSP when obtaining samples of the factor loadings and the error variances, and with the other approaches ROTATE, HT, SCAD, and LW, when obtaining a point estimate of the covariance matrix. In both cases, FABLE is faster than its competitors. When obtaining samples, FABLE is about 150200150-200 times faster than MGSP, with this ratio remaining fairly constant as $p$ increases. When obtaining a point estimator of the covariance, FABLE is at least four times as fast as ROTATE. Furthermore, obtaining posterior samples with FABLE is around $2.5$ times slower than obtaining the FABLE-point estimate, when averaged across replicates and dimensions.
Figure 1: Plots comparing the runtimes in seconds (in $\log_{10}$ scale) of FABLE with the competitors. The left panel compares FABLE with MGSP when obtaining posterior samples. The right panel compares FABLE, ROTATE, HT, SCAD, and LW, when obtaining a point estimate. Solid lines indicate average runtimes across replicates while shaded areas indicate minimum and maximum runtimes over replicates.
@@ 5.1 Setup
We used FABLE to identify associations between gene expressions in a high-throughput sequencing immunocellular dataset. Identifying gene expression associations in varying immune cell populations and between multiple lineages is critical to understanding the inflammatory response to infectious and chronic diseases (Ota et al.,, 2021; Saini et al.,, 2022). We consider the GSE109125 dataset, which contains RNAseq data from 127 highly purified immune cell lineages, including adaptive and innate lymphocytes, myeloid cells, mast cells, and neutrophils. These data were profiled using the ImmGen ULI pipeline (Yoshida et al.,, 2019).
@@
We first log-transform raw data counts $R_{ij}$ for the $i$th cell and $j$th gene as $y^{*}_{ij}=\log_{2}(R_{ij}+1)$ and center the data. We then use the genefilter package in R (Gentleman et al.,, 2023) to filter relevant genes before fitting the latent factor model. In this case, we filtered genes according to their variance and considered the top $10\%$ of genes with the highest variances. There were no cells with missing entries. After carrying out appropriate pre-processing of the dataset, we obtain expression data $\mathbf{Y}\in\mathbb{R}^{n\times p}$ for $p=5300$ genes measured on $n=205$ cells.
@@
We compare results from applying the three methods FABLE, MGSP, and ROTATE on the data. All three approaches fit the Bayesian latent factor model as described in Section 2. The estimated rank when using the $\mbox{JIC}(k)$ criterion as highlighted in Section 2.4.1 is $\widehat{k}=30$. For the other approaches MGSP and ROTATE, $k$ is estimated as described in Section 4.1. As a posterior predictive check, we investigated the average coverage of entrywise predictive intervals across all the entries of $\mathbf{Y}$. When using FABLE, the average predictive coverage is $0.97$. To further investigate model fit, we also looked into the proportion of variance explained for each gene using FABLE. The average proportion of variance explained across variables when using FABLE is $83\%$, indicating that the factor model is a good fit. We found the Gaussian assumption on the entries of $\mathbf{Y}$ well supported. However, a natural direction for future research is to extend the FABLE methodology to directly model count data, such as the $R_{ij}$s themselves.
5.2 Illustrating Blessing of Dimensions
We first highlight the blessing of dimensionality when implementing FABLE through the task of covariance submatrix estimation. We denote the full covariance matrix by $\Psi\in\mathbb{R}^{p\times p}$ where $p=5300$ and the covariance of any subset $\mathcal{I}\subset\{1,2,\ldots,p\}$ of the variables by $\Psi_{\mathcal{I}}\in\mathbb{R}^{|\mathcal{I}|\times|\mathcal{I}|}$. That is, $\Psi_{\mathcal{I}}$ is the submatrix of $\Psi$ corresponding to the indices in $\mathcal{I}$. Without loss of generality, we assume that the variable indices are in descending order of their variance; that is, variable $1$ has the largest variance and variable $p$ has the smallest variance. Our objective is to estimate the $100\times 100$ covariance submatrix corresponding to the $100$ genes with the highest variability. Suppose the indices of these genes are $\mathcal{I}_{0}=\{1,2,\ldots,100\}$. To estimate $\Psi_{\mathcal{I}_{0}}$ with FABLE, ROTATE, or MGSP, we could adopt two schemes, described as follows:
  1. (i)
    Consider gene expression data for variables with indices $\mathcal{I}_{0}$ and estimate $\Psi_{\mathcal{I}_{0}}$ only based on this data.
  2. (ii)
    For a given $p_{S}\geq 1$, let $\mathcal{A}(p_{S})=\{|\mathcal{I}_{0}|+1,\ldots,|\mathcal{I}_{0}|+p_{S}\}=\{100+1,\ldots,100+p_{S}\}$ denote the set of indices of genes with the next $p_{S}$ highest variances. We estimate $\Psi_{\mathcal{I}_{0}\cup\mathcal{A}(p_{S})}$ based on data for the variables with indices $\mathcal{I}_{0}\cup\mathcal{A}(p_{S})=\{1,\ldots,100+p_{S}\}$, and extract the submatrix $\Psi_{\mathcal{I}_{0}}$ from $\Psi_{\mathcal{I}_{0}\cup\mathcal{A}(p_{S})}$.
Scheme (i) is simply a special case of scheme (ii) with $p_{S}=0$ and $\mathcal{A}(0)=\phi$. We vary $p_{S}\in\{0\}\,\cup\,\{100j\,:\,j=1,\ldots,10\}\,\cup\,\{2000,4000\}$ and extract the relevant estimate of the covariance between the genes with indices in $\mathcal{I}_{0}$, namely $\widehat{\Psi}_{\mathcal{I}_{0}}(p_{S})$, for each $p_{S}$.
To obtain $\widehat{\Psi}_{\mathcal{I}_{0}}(p_{S})$ for a particular approach, we perform a random train-test split on $\mathbf{Y}$ with $50$ test samples and n50=155n-50=155 training samples, and train our model on the $155\times(100+p_{S})$ submatrix of $\mathbf{Y}$ as described earlier. We evaluate the performance of the estimate $\widehat{\Psi}_{\mathcal{I}_{0}}(p_{S})$ with the out-of-sample log-likelihood (OOSLL) of the $50\times 100$ test data corresponding to the indices $\mathcal{I}_{0}$. We carry out this analysis for $R=10$ independent replicates of the train-test split, and compare the methods FABLE, MGSP, and ROTATE on the basis of average OOSLL across replicates. Fitting the ROTATE approach by gradually increasing lambda0 as described in Section 4.1 did not perform well for this application. Instead, we obtain superior OOSLL by starting with lambda0 = 1 and gradually increasing it to lambda0 = 5 and lambda0 = 10, and provide the results for all three choices. The results are illustrated in Figure 2, with the OOSLL values available in Section H of the Supplementary Material.
@@ 5.2 Illustrating Blessing of Dimensions
Figure 2: Figure showing log-likelihood of the test set (y-axis) averaged over $R=10$ replicates of the train-test split when using additional genes (x-axis) to estimate the covariance of the relevant set of $100$ variables, with FABLE, MGSP, and ROTATE. The training set has $155$ cells and the test set has $50$ cells for each of the $R=10$ train-test split replicates.
From Figure 2, it is clear that both FABLE and MGSP exhibit superior performance to ROTATE. MGSP performs the best for smaller values of $p_{S}$, with both MGSP and ROTATE exhibiting an essentially monotonically decreasing trend as extra genes are added, so that including additional genes does not improve inferences on the genes of interest by using these methods. In sharp contrast, FABLE shows a rapid initial improvement in performance as extra genes are added, exhibiting a clear blessing of dimensionality. After a few hundred genes, the gain levels off, and eventually there is a modest decline, perhaps due to the need to add additional factors not related to the genes of interest when large numbers of additional genes are added. Thus, FABLE is the only approach among the three that benefits from considering additional variables when training the model, thus highlighting a “blessing of dimensionality” phenomenon.
The proposed method thus shows promising results and compares favourably with two state-of-the-art approaches to Bayesian factor analysis. The results obtained from FABLE come at a fraction of the computational budget compared to MGSP due to its embarrassingly parallel sampling scheme requiring no MCMC. When fitted on the full dataset on an M1 MacBook Pro with $32$ GB of RAM, FABLE only took $1.1$ seconds to compute the posterior mean while running MGSP with $3000$ MCMC and $1000$ burn-in iterates took $\sim 27$ minutes. This translates to a speed-up of close to $1600$ times for FABLE.
5.3 Train-Test Split
We also carry out a more straightforward train-test split exercise for the gene expression data application without subsetting the number of variables. For computational convenience, we filter the top $4\%$ genes with the highest variability from the raw data, obtaining a data matrix $\mathbf{Y}^{*}$ with $n=205$ as before and $p=2120$. We consider a sequence of $5$ training sample sizes $n_{T}\in\{110,130,150,170,190\}$. For each such $n_{T}$, we fit FABLE, MGSP, and the three variants of ROTATE with lambda0 in $\{1,5,10\}$ on the training set, obtain corresponding covariance matrix estimates, and then obtain the OOSLL of the $n-n_{T}$ held-out test samples with this estimate. We carry out this analysis for $R=10$ independent replicates of the train-test split and report average OOSLL for each $n_{T}$. The results are available in Figure 3, with the OOSLL values available in Section H of the Supplementary Material. The MGSP approach performed the best, followed by FABLE, with both outperforming all variants of the ROTATE approach with different values of the tuning parameter lambda0. All approaches show a naturally increasing trend for the OOSLL as the training sample size increases and the test sample size decreases.
Figure 3: Out-of-sample log-likelihood (y-axis) averaged over $R=10$ replicates for the direct train-test split exercise with gene expression data, using the FABLE, MGSP, and ROTATE approaches. The training sample size is varied from $110$ to $190$ in increments of $20$.
@@ 6 Discussion
In this paper, we develop a computationally scalable approach to fit high-dimensional Bayesian latent factor models with Gaussian data. Using a blessing of dimensions, the proposed approach bypasses the need for MCMC to provide FABLE-posterior samples of the covariance matrix for inference. Due to its embarrassingly parallel nature and reliance on independent samples, FABLE has immense computational benefits over current approaches dependent on MCMC. The FABLE posterior enjoys desirable theoretical properties, such as consistency and asymptotically accurate UQ of credible intervals.
This work has already led to several important extensions. Mauri and Dunson, 2025a extend the FABLE methodology to propose a scalable approach for fitting Bayesian multivariate logistic factor models. Their work is motivated by applications to joint species distribution modeling in ecology. The number of dimensions (species) for these problems is on the order of $p\sim 10^{4}-10^{5}$, rendering traditional MCMC-based approaches relying on data augmentation Gibbs samplers infeasible (Albert and Chib,, 1993; Polson et al.,, 2013). Other important extensions of FABLE include inference on covariance structure in high-dimensional multi-view data (Mauri and Dunson, 2025b, ) and multi-study factor analysis (Mauri et al.,, 2025), with compelling applications in high-throughput multi-omics analyses.
@@
Additional important modeling directions are to modify the FABLE inferential framework to allow more intricate hierarchical modeling. For example in Bayesian factor analysis, it is often of interest to (1) include covariate effects, (2) use more elaborate priors on the loadings to favor sparsity and adaptive selection of the number of factors, (3) consider more flexible latent factor distributions, and/or (4) allow nonlinear measurement structures. In exploratory analyses, we have observed good performance for ad hoc FABLE modifications by including covariates and for simulated data having non-Gaussian latent factor distributions. Problem (2) can potentially be addressed in the second stage via MCMC algorithms implemented in parallel for the different sparse regressions after inferring the latent factors using a sparse SVD in the first stage. Problem (4) may necessitate non-linear dimensionality reduction in the first stage. Obtaining concrete implementations with theoretical support for problems (1)-(4) is of future interest.
The FABLE approach crucially relies on the number of dimensions increasing to infinity as the number of samples grows. This is essential to ensure the accurate pre-estimation of the unobservable latent factors before leveraging this estimate to obtain exact Monte Carlo samples in an embarrassingly parallel manner. For smaller numbers of samples and dimensions where such an assumption may not hold, we conjecture that instead of a point pre-estimate of the unknown latent factors, they can be sampled from a distribution with probabilistic uncertainty around this pre-estimate, up to possible rotational constraints. As an extension of FABLE, we can then sample the latent factors, factor loadings, and error variances independently of one another from approximate posteriors while maintaining an embarrassingly parallel framework. Just like the proposed FABLE procedure, this modification would also entirely bypass MCMC to obtain posterior samples for the covariance matrix. We are actively working on this extension and have already obtained some promising results.
@@
The current work also hints at further investigation into the theoretical properties of Bayesian factor models. Although there are existing results focusing on posterior consistency with carefully chosen shrinkage priors on the factor loadings, these are primarily in the case where the idiosyncratic model errors are Gaussian. It would thus be interesting to explore how the blessing of dimensions influences both posterior consistency and uncertainty quantification in different settings, such as when the data are non-Gaussian. Furthermore, there remains a clear need to extend existing theoretical results on Bayesian factor models to encompass uncertainty quantification of the covariance matrix.